OpenStudy (anonymous):

find dy/dx using implicit differentiation: xe^y-ye^x=5

OpenStudy (john_es):

The implicit differentiation formula is, \[F(x,y)=xe^y-ye^x-5\Rightarrow \frac{dy}{dx}=-\frac{\partial F/\partial x}{\partial F/\partial y}\] Could you continue from this point?

OpenStudy (anonymous):

I'm still pretty confused, could you go furthur?

OpenStudy (anonymous):

further*

OpenStudy (abb0t):

I don't think he is referring to partial derivatives here, @John_ES

OpenStudy (abb0t):

Use product rule for xe\(^y\) and also with ye\(^x\)

OpenStudy (abb0t):

It is merely derivatives and tedious algebra solving for \(\sf \color{purple}{\frac{dy}{dx}}\)

OpenStudy (john_es):

Oh, well. It's true, there's another way to do the problem. I don't ask if @dn80919 has seen something about partial derivatives. Do you see them? If not, better follow the method @abb0t is explaining ;).

OpenStudy (anonymous):

hmmm..how would you take derivative of xe^y?

OpenStudy (abb0t):

Product rule: \(\sf \color{blue}{\frac{d}{dx}f(x)g(x)}\) = \(\sf \color{red}{ f(x)\frac{d}{dx}[g(x)]+g(x)\frac{d}{dx}[f(x)]}\)

OpenStudy (abb0t):

You should get: \(\sf \color{orange}{\large xe^y~\frac{dy}{dx}+e^y}\)

OpenStudy (john_es):

Are you sure of that? I think he should apply differentiation and not derivation. I mean, \[dx\cdot e^y+xe^ydy-dy\cdot e^x-ye^xdx=0\] And extracto common factor, \[dx(e^y-ye^x)+dy(xe^y-e^x)=0\]So, \[\frac{dy}{dx}=-\frac{e^y-ye^x}{xe^y-e^x}\]

OpenStudy (zzr0ck3r):

differentiation is the process of deriving the derivative to say "derive" means nothing unless you say what it is that you are deriving.

OpenStudy (abb0t):

@zzr0ck3r congrats on becoming green :)

OpenStudy (zzr0ck3r):

ty sir:)